Problem 8 a. y + 4y - sin (2 t) + + -1 b. 4y - 4y + y = 164/2 8. Evaluate the triple integral SSS w 2x DV, where W is the solid in th
Answer #1


SSS 23 2xdv bod region by Surface Z=22+42 + z = 2(x2+4°) Z=1 Z=1 z soon! SSS 2= x² +42 SS 1 - (x²+42) dxdy Х Osim projection